The sum of two numbers is 89. Three times the smaller number is 71 more than the larger. Find the numbers.
Answer by stanbon(75887)

The sum of two numbers is 89. Three times the smaller number is 71 more than the larger. Find the numbers.
-------
a + b = 89
3a = b+71
-------
a + b = 89
3a- b = 71
---------
4a = 160
a = 40
b = 89-40 = 49
----------------
Cheers,
Stan H.
-----------